OpenStudy (anonymous):

can someone please help me , i've been on here forevere trying to get help,but no one wants to help

OpenStudy (anonymous):

how do i graph 200x+100y<=700 x+y>=5 x>=1 x>=0

jimthompson5910 (jim_thompson5910):

graph each one at a time to graph 200x+100y<=700, you 1) graph the line 200x+100y=700 2) make it a solid line 3) shade below the solid line

jimthompson5910 (jim_thompson5910):

graphing x+y>=5 1) graph the line x+y=5 2) make it a solid line 3) shade above the solid line

jimthompson5910 (jim_thompson5910):

graphing x >= 1 1) graph the vertical line at x = 1 2) make it a solid line 3) shade to the right of this vertical solid line

jimthompson5910 (jim_thompson5910):

graphing y >= 0 (assuming you meant to say y >= 0 instead of x >= 0) 1) graph the horizontal line at y = 0 2) make it a solid line 3) shade above this horizontal solid line

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

after you've graphed each individual inequality (on the same xy axis) you then shade in the region that's common to all 4 inequalities...so basically where all 4 inequalities overlap this will be the solution region

OpenStudy (anonymous):

alright